4. Gravetter/Wallnau/Forzano, Essentials Chapter 1- End-of-chapter question 18 For the following scores, find the value of each expression 4 x2 =
Answer #1

1. X=2+2+4+2=11

2. X2-32 + 22 + 42 + 22 = 9 + 4 + 16 + 4 = 33

3. X-2=11-2=9

4. 〉 (X-2) (3-2) + (2-2):4-2) + (2-2) = 1+0+2+0=3